OpenStudy (anonymous):

NO IDEA where to start :/ When finding the sum of the arithmetic series 4 + 1 + –2 + –5 + –8 + –11, you will have a step of work that contains the number –3. Use complete sentences to describe how the –3 was calculated and what it means.

OpenStudy (cwrw238):

notice 1-4 = -3 and -2-1 = -3 in fact the next term of the series is obtained by adding -3 to current term this defines an arithmetic series. this is called the common difference (d) of the series. the series can be written as a, a+d, a + 2d, a+ 4d ......... where a = first term

OpenStudy (anonymous):

Thank you!

OpenStudy (cwrw238):

yw